OpenStudy (anonymous):

Find the derivative of the function. h(x) = log6(x2 + x)

terenzreignz (terenzreignz):

Seems like a true-blue application of the chain rule?

OpenStudy (anonymous):

that, and the fact that \[\log_6(x)=\frac{\ln(x)}{\ln(6)}\] so the derivative of \[\log_6(x)\] is \[\frac{1}{\ln(6)x}\]

terenzreignz (terenzreignz):

anytime you're ready @arshia93 ^_^

OpenStudy (anonymous):

i put \[\frac{ 1 }{ (2x+1)\ln6 }\] but that came out incorrect.

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Chain rule though. Bit by bit...

terenzreignz (terenzreignz):

Remember...\[\Large \frac{d}{dx}f(g(x))= f'(g(x)) g'(x)\]

terenzreignz (terenzreignz):

So you could take \[\Large f(x) = \log_6(x)\]\[\Large g(x) = x^2 +x\]

OpenStudy (anonymous):

im supposed to fin detivative of x^2+x correct? So i included (2x+1) with the fraction from log6

OpenStudy (anonymous):

is it not included in the fraction?

terenzreignz (terenzreignz):

Let me be more blunt about it ^_^ \[\Large \frac{d}{dx} \log_6[g(x)]= \frac{g'(x)}{g(x)\ln (6)}\] PLEASE review the chain rule.

OpenStudy (anonymous):

\[\frac{ 2x+1 }{ \ln6(x^2+x) }?\]

OpenStudy (anonymous):

when you use the chain rule never change the input!!

OpenStudy (anonymous):

yes